OpenStudy (anonymous):

Classify the system and determine the number of solutions. (Check all that apply.) 6x + 3y = -12 2x + y = -6 no solution consistent infinitely many solutions inconsistent

OpenStudy (anonymous):

do you know this? Use substitution to solve the system of equations. y + 1 = x -2x + 3y = 2 (1/4, 1) (5, 4) (-6, 3/2) (1/2, 15)

OpenStudy (anonymous):

Just no solution?

OpenStudy (anonymous):

thank you! Do you know how to do the other?

OpenStudy (anonymous):

thanks!!